# Cloud gap-filling with deep learning for improved grassland monitoring

### v1

@@ 5.3 Recovery of masked events
In Section 5.2, we assessed the impact of the SF algorithm on the two event detection methods. This section investigates cases where grassland parcel activities are unobservable through optical data (due to consecutive cloudy observations), rendering NDVI-based detection inadequate for identifying such events. For this purpose, we selected 359 cases from the dataset with a single mowing event in order to intentionally hide them. On these cases, we artificially mask out the NDVI value from the day before the event and the following values that could reveal grassland activity (e.g., lower NDVI values). This leads to gaps of 3-7 consecutive missing values (24-48 days).
In Table 6, we compare the performance of each mowing detection algorithm applied to i) the raw NDVI series (no interpolation) and ii) the reconstructed NDVI time series using the SF model. Both algorithms perform poorly without interpolation due to lack of optical information. On the other hand, SF is able to recover nearly half of the actual events for MDA I and over half for MDA II, with relatively high precision values of 0.63 and 0.58, respectively. This demonstrates the robustness of SF in improving detection accuracy and reconstructing NDVI drops, even for large temporal gaps, making it a promising approach when optical data fails to capture event information.
@@
Algorithm Interpolation Recall Precision F1-score MDA I - 0.036 0.074 0.048 SF 0.521 0.810 0.634 MDA II - 0.123 0.128 0.134 SF 0.655 0.701 0.677
In Fig. 13, we illustrate how the accuracy of event detection is affected by the time-window (tolerance) within which we consider an event as correct. The SF method reaches its optimal recall performance of around 0.5 for MDA I and 0.6 for MDA II after 8 days, which remains relatively stable for higher tolerance parameter as we are reaching the default threshold (i.e., 12 days). Interestingly, even for very low tolerance values, it achieves satisfactory performance, especially considering cases of completely hidden events. In contrast, non-interpolated NDVI time series result in extremely low recall values, irrespective of the tolerance value.
@@
Figure 14: Spatial visualization of NDVI time series for a parcel. (a) Actual NDVI extracted from the Sentinel-2 cloud-free acquisition dates, (b) artificial NDVI time series created after masking the values on the dates during and after the observed mowing event, which are highlighted in red (from 21/06 to 09/07)). For further visualization, refer in Fig. S8 of A.
5.4 DNN mowing detection architecture
@@
In this section, we utilized the SF architecture to detect mowing events (see Section 4.4) and compared its performance with the algorithms mentioned earlier. To evaluate the model, we conducted a 3-fold cross-validation using the photo-interpretation dataset (refer to Section 3.2.2). The dataset was divided into three folds at the parcel level, taking into account the date of the annotated mowing events to ensure fairness. Specifically, each fold includes parcels with a distribution of event dates that matches that of the entire dataset. However, even though we split on the parcel level, the training of the models performed at the pixel level. This approach helped us obtain reliable results for evaluating the SF architecture. Finally, for our suggested approach, which requires 6-day gap time series, we trained two models; i) in the one we used time series interpolated with AI and ii) on the other we used the reconstructed time series using the SF model. These results were compared with the ones obtained for each MDA after applying i) the raw time series without any interpolation, ii) the same time series after Akima interpolation, and the subsequent acquisition of a 6-day temporal resolution, and iii) the corresponding results after implementing the SF model.
Table 7 displays the results on the mowing event detection task. MDA I achieved the highest recall (0.875), albeit with a relatively low precision of (0.686), resulting in suboptimal overall performance. Notably, Akima interpolation significantly improved precision without sacrificing recall. A similar trend was observed for MDA II, with a slight decrease in overall recall, leading to a substantial increase in precision. In both cases, the use of SF-derived time series resulted in significantly improved precision while maintaining similar recall levels, leading to an increased overall f1-score of approximately 0.83.
@@
Our suggested approach outperforms threshold-based methods in identifying mowing events, especially with SF-processed NDVI input, achieving optimal results with significantly higher precision (0.83 and 0.86). In both cases of our proposed implementation, the mean f1-score exceeded 80%, indicating highly effective performance in mowing event detection. Notably, our suggested supervised approach achieved f1-score close to 85% when incorporating SF for interpolation, representing an approximately 2% increase over the Akima interpolation case. This is a significant improvement, especially given that we have used very dense NDVI time series (i.e., minimal cloud coverage), making the majority of events relatively easy to detect for both input scenarios.
@@ 6.1 Cloud filtering for event detection
Cloud cover can be so extensive at times that it results in gaps in the time series that may span more than a month. In Section 5.3, we illustrated how the SF algorithm produces a continuous time series, allowing for the identification of mowing events that would have been missed if only optical imagery had been used. However, cloud masks may not always detect the presence of clouds accurately, resulting in false measurements and inaccurate event detection. According to hardy2021sen2grass , there is a trade-off between data availability and accuracy when setting the cloud cover threshold for grassland monitoring systems, which depends on the end-user’s preferences.
Apart from reconstructing missing information, the SF methodology also serves as a supplementary cloud filtering mechanism. By incorporating SAR and optical data, we observe that it can identify and disregard available Sentinel-2 measurements that result in sudden drops of NDVI, when no correlation between the two components is justified.
In Fig 15, we illustrate a typical scenario where the sen2Cor algorithm fails to detect high cirrus clouds (mid-April), resulting in the inclusion of a noisy measurement that could potentially be misinterpreted as an event from the mowing event algorithms. As shown, the SF methodology ignores the NDVI drop, even though it was included into the algorithm’s input data and instead, it generates an output following the general upward trend.
Figure 15: Example of a case where the SF algorithm acts as a cloud filtering mechanism. The top three RGB images have been derived from consecutive Sentinel-2 captures. The first and the last are clear-sky images while the middle one suffers from high cirrus clouds. In the plot below are presented the actual NDVI time series (orange) and the predicted ones using the SF algorithm (green). The blue, red and purple vertical dotted lines refer to the dates that the aforementioned images were captured.
Taking the aforementioned into account, a filtering routine can be developed. By employing a particular threshold difference between the actual NDVI value and the one generated by the SF model, timestamps that are potentially affected by clouds can be identified and excluded. This additional technique can improve the initial cloud masking process, resulting in more accurate time series data for monitoring agricultural land and making it more suitable for the subsequent tasks.
@@ 6.2 Gap-filling ablation study
To determine the individual contributions of different components within the SF model, we conducted an ablation study. This study specifically analyzed the NDVI gap-filling performance across six varied input data configurations. The baseline scenario includes all the features described earlier in Table 2, while the other five scenarios are constructed by combining specific features, and their derived products. These scenarios involve combinations or single deployments of the $\mathrm{\sigma_{0}}$, coherence, and NDVI features to showcase various input configurations. All scenarios maintain consistent model parameter configurations and use the same training and testing pixels from the datasets (see Table 4).
Table 8 presents the MAE and $\mathrm{R^{2}}$ of the different scenarios. The performance of the model without using as input the cloud-free NDVI time series is substantially lower. In these cases, the average MAE was more than 0.051, and the $\mathrm{R^{2}}$ remained below 0.79. The incorporation of the available NDVI measurements leads to a significant improvement, with a decrease of the average MAE by 26-29%, depending on the case, and an increase in $\mathrm{R^{2}}$ by approximately 0.09, which corresponds to an increase of around 11%. Specifically, MAE gets average values below 0.04, and $\mathrm{R^{2}}$ over 0.87. Finally, the best results were obtained with the inclusion of all derived features, as expected.
@@
Input Features MAE $\mathrm{R^{2}}$$\mathrm{coherence}$} 0.0533 $\pm$ 0.007 0.785 $\mathrm{\sigma_{0}}$} 0.0529 $\pm$ 0.006 0.787 $\mathrm{coherence,\sigma_{0}}$} 0.0518 $\pm$ 0.006 0.789 $\mathrm{NDVI,coherence}$} 0.0395 $\pm$ 0.004 0.876 $\mathrm{NDVI,\sigma_{0}}$} 0.0389 $\pm$ 0.004 0.876 $\mathrm{NDVI,coherence,\sigma_{0}}$} 0.0370 $\pm$ 0.004 0.881
@@ 6.3 Spatial generalization
To gain additional insights, we conducted an experiment to assess the model’s behavior when trained on a limited number of regions. This allowed us to evaluate the model’s ability to generalize spatially to regions that were not included in the original training process. Additionally, by examining how the number of training samples affects the quality of the training process, we can determine the minimum number required to achieve satisfactory results.
The optimal generalization capability of the SF model depends heavily on the size of the training dataset. To accurately learn from the distribution of missing values and fill NDVI gaps, it is crucial to have a representative sample of training instances across the entire Lithuania. Fig. 16 demonstrates the impact of increasing the number of training instances and the regions used for training on MAE. The darker colors indicate more training regions used. While including more training samples leads to a decrease in MAE, it appears that using training samples from different regions is more crucial. Interestingly, by utilizing only three regions, we can achieve near-optimal results. Despite the limited number of training regions, in many cases the model achieves near-optimal performance. This may be due to the uniform mowing policy and similar grassland management practices throughout Lithuania, resulting in similar grassland vegetation profiles. It can be also attributed to the fact that there are some regions that are larger (e.g., Region 1 and 2), which include diverse cases, and thus when they are used for training, we can more accurately represent the overall distribution. However, it’s essential to consider the cloud coverage distribution of each training region, as this can affect the model’s generalization ability beyond those regions. If the cloud coverage profile of the training samples differs significantly from the overall distribution, the model may overfit to the training regions and fail to generalize well. For more information, in the A (Table S4) we present the MAE computed on the inference subset of each region, and overall, when the model is trained on one, two, or three regions.
Figure 16: Scatter plot comparing the MAE with different number of training pixels. The color of each point indicates the number of training regions, as shown in the right color bar.
@@ 6.4 Limitations and future work
In recent years, numerous studies have concentrated on employing neural networks to combine SAR and optical data to tackle the problem of discontinuous optical image time series. However, none of these studies have investigated incorporating the available optical data directly into the model’s architecture. According to garioud2021 , such an approach could potentially cause the model to ignore the remaining input features because the target and input data would be identical at several timestamps. Our observations during the training of SF using NDVI time series as an additional input feature confirmed this issue, as we noticed a high correlation between the input and output NDVI time series. To address this concern, we used temporal weighting for each input sample, as mentioned in Section 4.3.3. We assigned masked NDVI values a weight of 0.75, while non-masked values received a weight of 0.25. Moreover, we assigned a weight of 0 to any value corresponding to an interpolated timestamp, as we used interpolation to produce a complete time series.
@@
Our selection of the CNN-RNN architecture over more advanced DNN architectures (e.g., Transformers) is motivated by the fact that it has been already suggested and applied in a similar research task (zhao2020, ) and it also offers a computational efficient approach. Furthermore, our primary objective was to emphasize over the contribution of integrating the Sentinel-2 measurements into the model’s feature space. Overall, we are confident that this architecture adequately presents our goal of improving the reconstruction of missing NDVI values and enhancing mowing detection accuracy. We also intend to explore a broader range of architectures in future research.
@@
By leveraging the ability of neural networks to learn complex temporal patterns, we capture grassland phenology information and dependencies. NDVI generally requires approximately 10 days to recover half of the value lost after a grassland mowing event (kolecka2018, ). In Lithuania, it is uncommon to find cases where no optical cloud-free images has been captured during such long periods, especially during the summer months. The majority of these gaps occur before June, whereas most mowing events take place from early June onward. Fig. 17 depicts the temporal distribution of extended gaps. Extended gaps refer to occasions where pixels have been masked out because of cloud coverage for at least the three previous Sentinel-2 acquisitions. This duration is considered sufficient to conceal a mowing event. Consequently, while the model can effectively uncover hidden mowing events, it is important to note that, in our study, such occurrences are anticipated to be relatively infrequent (5-10%).
Figure 17: Frequency of extended gaps on each date. Extended gaps are considered cases of pixels in a Sentinel-2 image that have been masked out due to cloud coverage for at least the three previous acquisitions.
@@
In Section 6.3, we showcased the generalization ability of our model across Lithuania. Our future goal is to extend this approach to other European countries with diverse grassland mowing and activity policies, as well as varying cloud coverage distributions. This would lead in the creation of a pan-European model that could be highly valuable for CAP monitoring.
@@
Moreover, our approach can be adapted through minor modifications to reconstruct other remote sensing indices, such as Soil Adjusted Vegetation Index (SAVI), and Normalized Burn Ratio (NBR), or even raw optical spectral bands. These indices are particularly beneficial for monitoring vegetation or detecting farming practices, including tillage and stubble burning. Consequently, a more versatile methodology can be developed to generate one or multiple relevant indices, depending on the task. Furthermore, monitoring various land cover vegetation types is possible, including forests, arable croplands (such as wheat, maize, or barley) and orchards.

### v2

@@ 5.3 Recovery of masked events
In Section 5.2, we assessed the impact of the SF algorithm on the two event detection methods. This section investigates cases where grassland parcel activities are unobservable through optical data (due to consecutive cloudy observations), rendering NDVI-based detection inadequate for identifying such events. For this purpose, we selected 359 cases from the dataset with a single mowing event in order to intentionally hide them. In these cases, we artificially masked out the NDVI value from the day before the event and the following values that could reveal grassland activity (e.g., lower NDVI values). This led to gaps of 3-7 consecutive missing values (24-48 days).
In Table 6, we compare the performance of each mowing detection algorithm applied to i) the raw NDVI series (no interpolation) and ii) the reconstructed NDVI time series using the SF model. Both algorithms performed poorly without interpolation due to a lack of optical information during the events. On the other hand, SF was able to recover nearly half of the actual events for MDA I and over half for MDA II, with relatively high F1-score values of 0.63 and 0.68, respectively. This demonstrates the reliability of SF in improving detection accuracy and reconstructing NDVI drops, even for large temporal gaps, making it a promising approach when optical data fails to capture event information. Additionally, in Fig. S10 and S11 (refer to section A), spatial and temporal distribution for the recall is illustrated respectively, highlighting the robustness of the results.
@@
Algorithm Interpolation Recall Precision F1-Score MDA I - 0.036 0.074 0.048 SF 0.521 0.810 0.634 MDA II - 0.123 0.148 0.134 SF 0.655 0.701 0.677
In Fig. 14, we illustrate how the accuracy of event detection is influenced by the time-window (tolerance) used to define an event as correct. The SF method reached its optimal recall performance of around 0.5 for MDA I and 0.6 for MDA II after 8 days, which remained relatively stable for higher tolerance parameters as we were reaching the default threshold (i.e., 12 days). Interestingly, even for very low tolerance values, it achieved satisfactory performance, especially considering cases of completely hidden events. In contrast, non-interpolated NDVI time series resulted in extremely low recall values, irrespective of the tolerance value.
@@
Figure 15: Spatial visualization of NDVI time series for a parcel. (a) Actual NDVI extracted from the Sentinel-2 cloud-free acquisition dates, (b) artificial NDVI time series created after masking the values on the dates during and after the observed mowing event, which is highlighted in red (from 21/06 to 09/07)). For further visualization, refer to Fig. S12 of A.
5.4 Deep learning mowing detection algorithm
@@
In this section, we utilized the SF architecture to detect mowing events (see Section 4.4) and compared its performance with the algorithms mentioned earlier. To evaluate the model, we conducted a 3-fold cross-validation using the photo-interpretation dataset (refer to Section 3.2.2). The dataset was divided into three folds at the parcel level, taking into account the date of the annotated mowing events to guarantee fairness. Specifically, each fold contained cases with event date distributions that mirrored those of the entire dataset. However, even though we split on the parcel level, the training of the models was performed at the pixel level. This approach helped us obtain reliable results for evaluating the SF architecture. Finally, for our proposed approach, which required a 6-day interval time series, we trained two models: i) one using time series interpolated with the Akima method, and ii) another using time series constructed using the SF model. These results were compared with those obtained from each MDA under the following conditions: i) using raw, unprocessed time series without interpolation, ii) using the same time series after Akima interpolation to achieve a 6-day temporal resolution, and iii) using the corresponding results after applying the SF model.
Table 7 displays the results of the mowing event detection task. MDA I achieved the highest recall (0.875), albeit with relatively low precision (0.686), resulting in suboptimal overall performance. Notably, Akima interpolation significantly improved precision without sacrificing recall. A similar trend was observed for MDA II, with a slight decrease in overall recall, leading to a substantial increase in precision. In both cases, the use of SF-derived time series resulted in significantly improved precision while maintaining similar recall rates, leading to an increased overall F1-score of approximately 0.83.
@@
Our suggested approach outperformed threshold-based methods in identifying mowing events, especially with SF-processed NDVI input, achieving optimal results with significantly higher precision (0.83 and 0.86). In both cases of our proposed implementation, the mean F1-score exceeded 80%, indicating highly effective performance in mowing event detection. Remarkably, our suggested supervised approach achieved F1-score close to 85% when incorporating SF for interpolation, which marks an improvement of approximately 2% over the Akima case, highlighting the added value of NDVI reconstruction. This represents a notable improvement, considering the utilization of highly dense NDVI time series (i.e., minimal cloud coverage), making the detection of the majority of events relatively easy for both input scenarios.
@@ 6.1 Cloud filtering for event detection
Cloud cover can be so extensive at times that it results in gaps in the time series that may span more than a month. In Section 5.3, we illustrated how the SF algorithm produces a continuous time series, allowing for the identification of mowing events that would have been missed if only optical imagery had been used. However, cloud masks may not always detect the presence of clouds accurately, resulting in false measurements and inaccurate event detection. According to [71], there is a trade-off between data availability and accuracy when setting the cloud cover threshold for grassland monitoring systems, which depends on the end-users preferences.
Apart from reconstructing missing information, the SF methodology also serves as a supplementary cloud filtering mechanism. By incorporating SAR and optical data, we observe that it can identify and disregard available Sentinel-2 measurements that result in sudden drops of NDVI when no correlation between the two components is justified.
In Fig 16, we illustrate a typical scenario where the sen2Cor algorithm fails to detect high cirrus clouds (mid-April), resulting in the inclusion of a noisy measurement that could potentially be misinterpreted as an event from the mowing event algorithms. As shown, the SF methodology ignores the NDVI drop, even though it was included in the algorithm’s input data and instead, it generates an output following the general upward trend.
Figure 16: Example of a case where the SF algorithm acts as a cloud filtering mechanism. The top three RGB images have been derived from consecutive Sentinel-2 captures. The first and the last are clear-sky images while the middle one suffers from high cirrus clouds. In the plot below are presented the actual NDVI time series (orange) and the predicted ones using the SF algorithm (green). The blue, red, and purple vertical dotted lines refer to the dates that the aforementioned images were captured.
Taking the aforementioned into account, a filtering routine can be designed. By employing a particular threshold difference between the actual NDVI value and the one generated by the SF model, timestamps that are potentially affected by clouds can be identified and excluded. This additional technique can improve the initial cloud masking process, resulting in more accurate time series data for monitoring agricultural land and making it more suitable for subsequent tasks.
@@ 6.2 Gap-filling ablation study
To determine the individual contributions of different components within the SF model, we conducted an ablation study. This study specifically analyzed the NDVI gap-filling performance across seven varied input data configurations. The baseline scenario includes all the features described earlier in Table 2, while the other six scenarios are constructed by combining specific features, and their derived products. These scenarios involve combinations or single deployments of the $\mathrm{\sigma_{0}}$ (similar to [29]), coherence, and NDVI features to showcase various input configurations. All scenarios maintained consistent model parameter configurations and used the same training and validation pixels from the datasets (see Table 4). Table 8 presents the MAE and $\mathrm{R^{2}}$ of the different scenarios. The performance of the model without utilizing the cloud-free NDVI time series as input is substantially lower. In these cases, the average MAE exceeds 0.051, and the $\mathrm{R^{2}}$ remains below 0.79. Incorporating SAR data with the available NDVI measurements led to a substantial improvement, reducing the average MAE by 26-29% depending on the scenario, and increasing $\mathrm{R^{2}}$ by approximately 0.09, corresponding to an enhancement of about 11%. Specifically, MAE gets average values below 0.04, and $\mathrm{R^{2}}$ over 0.87. Notably, even when using only the NDVI component, there was a significant enhancement in terms of MAE, with a decrease of 0.01 compared to using only SAR input features. This outcome could be justified, as the model learns from the statistical profile of regional NDVI, operating as a deep learning-oriented interpolation approach, unaffected by less-correlated SAR data. Consequently, the results are similar to or outperform traditional baseline interpolation methodologies, achieving the best MAE of 0.43 (see Section 5). Finally, the best results were obtained with the inclusion of all derived features, pointing out the contribution of SAR components to the model.
@@
Input Features MAE $\mathrm{R^{2}}$$\mathrm{coherence}$} 0.0533 $\pm$ 0.007 0.785 $\mathrm{\sigma_{0}}$} 0.0529 $\pm$ 0.006 0.787 $\mathrm{coherence,\sigma_{0}}$} 0.0518 $\pm$ 0.006 0.789 $\mathrm{NDVI}$} 0.0413 $\pm$ 0.008 0.841 $\mathrm{NDVI,coherence}$} 0.0395 $\pm$ 0.004 0.876 $\mathrm{NDVI,\sigma_{0}}$} 0.0389 $\pm$ 0.004 0.876 $\mathrm{NDVI,coherence,\sigma_{0}}$} 0.0364 $\pm$ 0.004 0.885
Additionally, to thoroughly understand the role of different elements in our CNN-RNN architecture, we conducted an ablation study by systematically removing or modifying parts of the network and evaluating the performance impact. This approach allowed us to determine the optimal hyperparameter setup. In Table S4 (see A), detailed results of the tested architecture variants are presented in terms of MAE and $\mathrm{R^{2}}$. Specifically, the contribution of the CNN layers at the beginning of the network is evident, as they significantly enhanced the overall performance. The number of parameters for each model configuration is also provided, emphasizing that our models are lightweight and suitable for deployment on conventional infrastructures, with no requirement for GPU acceleration.
Finally, in Table S5 (see A), we present the results of a grid search evaluating different training weight parameters ($\mathrm{w_{\alpha}}$, $\mathrm{w_{\beta}}$) for both masked and non-masked input NDVI values. The results indicate that smaller $\mathrm{w_{\beta}}$ values compared to $\mathrm{w_{\alpha}}$, led to optimal performance. Conversely, assigning equal weight parameters (i.e., setting both $\mathrm{w_{\alpha}}$ and $\mathrm{w_{\beta}}$ to 1) resulted in slightly poorer MAE, while worse results are obtained with larger $\mathrm{w_{\beta}}$ parameters. These findings underscore our initial statement to prioritize attention on masked timestamps to mitigate the NDVI leakage issue when used as both input and target values (see Section 4.3.3).
@@ 6.3 Spatial generalization
To gain additional insights, we conducted a regional cross-validation experiment to assess the model’s behavior when trained on a limited number of regions. This allowed us to evaluate the model’s ability to generalize spatially and transfer to regions that were not included in the original training process. Additionally, by examining how the number of training samples affects the quality of the training process, we can determine the minimum number required to achieve satisfactory results.
The optimal generalization capability of the SF model depends heavily on the size of the training dataset. To accurately learn from the distribution of missing values and fill NDVI gaps, it is crucial to have a representative sample of training instances across the entire Lithuania. Fig. 17 demonstrates how increasing the number of training instances and the regions used for training affects the MAE. The darker colors indicate more training regions used. While including more training samples led to a decrease in MAE, it appears that using training samples from different regions is more crucial. Interestingly, by utilizing only three regions, we can achieve near-optimal results.
Figure 17: Scatter plot comparing the MAE with different numbers of training pixels. The color of each point indicates the number of training regions, as shown in the right color bar.
Despite the limited number of training regions, in many cases, the model achieved near-optimal performance. This may be due to the uniform mowing policy and similar grassland management practices throughout Lithuania, resulting in similar grassland vegetation profiles. It can be also attributed to the fact that some regions are larger (e.g., Region 1 and 2), which include diverse cases, and thus when they are used for training, we can more accurately represent the overall distribution. However, it’s essential to consider the cloud coverage distribution of each training region, as this can affect the model’s transferability beyond those regions. If the cloud coverage profile of the training samples differs significantly from the overall distribution, the model may overfit to the training regions and fail to generalize well. For more information, in the A (Table S6) we present the MAE computed on the validation subset of each region, and overall, when the model was trained on one, two, or three regions.
Furthermore, to address concerns regarding the generalization capability of our model for the downstream task of mowing detection, we conducted an additional analysis of mowing detection performance across various MDA using the reconstructed NDVI produced from the trained models mentioned above. This analysis aimed to evaluate performance under different training region configurations, providing insights into the model’s ability to detect events across diverse geographical regions. Our results corroborate this argument, demonstrating lower average F1-scores and increased variability when the SF model is trained using data from only one region. As the number of training regions increases, performance improves, approaching F1-scores comparable to those achieved after optimal SF model implementation. Specifically, for MDA I and MDA II, the average F1-scores reached 0.82 and 0.83, respectively, with three training regions included. For detailed information, refer to Table S8 in A.
@@ 6.4 Limitations and future work
In recent years, numerous studies have concentrated on employing neural networks to combine SAR and optical data to tackle the problem of discontinuous optical image time series. However, none of these studies have investigated incorporating the available optical data directly into the model’s architecture. According to [28], such an approach could potentially cause the model to ignore the remaining input features because the target and input data would be identical at several timestamps. Our observations during the training of SF using NDVI time series as an additional input feature confirmed this issue, as we noticed a high correlation between the input and output NDVI time series. To address this concern, we used temporal weighting for each input sample, as mentioned in Section 4.3.3. We assigned masked NDVI values a weight of 0.75, while non-masked values received a weight of 0.25. Moreover, we assigned a weight of 0 to any value corresponding to an interpolated timestamp, as we used Akima interpolation to produce a complete time series. Additionally, an ablation study to determine the individual contributions of different components to the model showcases the clear benefits of including both optical and SAR information.
@@
Our selection of the CNN-RNN architecture over more advanced deep learning architectures (e.g., Transformers) is motivated by the fact that it has been already suggested and applied in a similar research task [29] and it also offers a computationally efficient approach. Furthermore, our primary objective was to emphasize the contribution of integrating the Sentinel-2 measurements into the model’s feature space. Overall, we are confident that this architecture adequately presents our goal of improving the reconstruction of missing NDVI values and enhancing mowing detection accuracy. We also intend to explore a broader range of architectures in future research.
@@
Utilizing neural networks’ ability to learn complex temporal patterns, we extracted grassland phenology information and dependencies. Typically, NDVI requires approximately 10 days to recover half of the value lost after a grassland mowing event [9]. In Lithuania, it is uncommon to find cases where no optical cloud-free images have been captured during such long periods, especially during the summer months. The majority of these gaps occur before June, whereas most mowing events take place from early June onwards. Fig. S13 in A depicts the temporal distribution of extended gaps. Extended gaps refer to occasions where pixels have been masked out because of cloud coverage for at least the three previous Sentinel-2 acquisitions. This duration is considered sufficient to conceal a mowing event. Consequently, while the model can effectively uncover hidden mowing events, it is important to note that such occurrences are anticipated to be relatively infrequent (5-10%). This is also evident by combing Fig. S6 and S7 on A.
Another challenge of the study relates to the photo-interpretation process employed for annotating mowing events. Despite efforts to ensure accuracy, human error remains inherent in manual interpretation tasks, potentially leading to inconsistencies or inaccuracies in event identification. Additionally, distinguishing between mowing and grazing events solely based on satellite imagery proved challenging. As a result, some events annotated as mowing may actually correspond to grazing activity, introducing a degree of ambiguity into the dataset.
@@
In Section 6.3, we showcased the generalization ability of our model across Lithuania. Our future goal is to extend this approach to other European countries with diverse grassland mowing and activity policies, as well as varying cloud coverage distributions. This would lead to the creation of a pan-European model that could be highly valuable for CAP monitoring.
@@
Moreover, our approach can be adapted through minor modifications to reconstruct other remote sensing indices, such as the Soil Adjusted Vegetation Index (SAVI), Normalized Burn Ratio (NBR), or even raw optical spectral bands. These indices are essentially beneficial for monitoring vegetation or detecting farming practices, including tillage and stubble burning. Consequently, a more versatile methodology can be developed to generate one or multiple relevant spectral indices, depending on the task. Furthermore, monitoring various land cover vegetation types is possible, including forests, arable croplands (such as wheat, maize, or barley), and orchards.